# Visual and Textual Prompts for Enhancing Emotion Recognition in Video

### v1

@@ IV-B Dataset details
IV-C Action units
TABLE IV: Comparison of state-of-the-art (SOTA) methods for zero-shot emotion recognition. Previous approaches utilize various types of visual prompts: C: Crop, O: Box, R: Blur Reverse, E: Circle, N: Number, and A: Action Units.
@@ IV-C Action units
IV-D Quantitative Results
@@
IV-E Visual Prompting
Table IV presents a comparative analysis of state-of-the-art (SOTA) methods for zero-shot emotion recognition, evaluating performance across Easy, Medium, and Hard difficulty tiers using Accuracy (Acc) and F1-score (F@1) metrics. It compare the effectiveness of different visual prompting strategies utilized by Qwen2-VL. The proposed SoVTP (Ours) method, which integrates numeric annotations (N), bounding boxes (O), and facial action units (A) as visual prompts, achieves superior performance, significantly outperforming existing approaches. On the Easy tier, SoVTP attains 70.59% Accuracy and 58.00% F@1, a substantial improvement over the second-best method, ReCLIP (35.29% Acc, 24.85% F@1). Similarly, in the Medium tier, SoVTP achieves 48.00% Accuracy and 44.94% F@1, surpassing RedCircle (36.00% Acc, 33.20% F@1) and ReCLIP (36.00% Acc, 32.29% F@1). For the challenging Hard tier, SoVTP maintains robustness with 38.57% Accuracy and 28.25% F@1, outperforming ReCLIP (34.29% Accuracy, 20.75% F@1) and RedCircle (31.43% Acc, 17.71% F@1). In total metrics, SoVTP achieves 45.54% Accuracy and 33.53% F@1, representing a 10.72% absolute Accuracy gain over ReCLIP (34.82% Accuracy, 23.13% F@1) and a 13.40% Accuracy gain over RedCircle (32.14% Accuracy, 19.33% F@1). Notably, the baseline method (Qwen2-VL) with plain text prompts lags significantly across all tiers (Total: 23.21% Accuracy, 10.06% F@1), underscoring the critical role of structured visual prompts. The comparatively lower performance of RedCircle and ReCLIP can be attributed to their reliance on cropping and blurring techniques. This approach increases the model’s focus on specific facial features but sacrifices essential body language and contextual cues, which are vital in more complex situations where emotional cues extend beyond the face. In contrast, the SoVTP approach retains a broader field of view, capturing body language and contextual information. This holistic approach accounts for the significant performance gains observed with SoVTP, especially in challenging conditions, where both facial and contextual cues are crucial for accurate emotion recognition.
TABLE V: Ablation study for vision and text prompts on Qwen2-VL. Baseline: represents the model’s performance without any additional prompts. Muscle: indicates a visual prompt that uses action units. Muscle+Context: adding extracted context to muscle prompts. Muscle+Context+Body: adding extracted body language to context and muscle prompts. Ours: adding extracted others’ emotions to muscle, context and body language prompts.
@@ IV-E Visual Prompting
IV-F Ablation Study
Table V presents an ablation study of vision and text prompt components on the Qwen2-VL framework, evaluating the incremental impact of integrating multi-modal cues for zero-shot emotion recognition. The baseline model, devoid of additional prompts, achieves minimal performance (Total: 23.21% Acc, 10.06% F@1), particularly struggling in Easy (11.76% Acc) and Medium (20.00% Acc) tiers. Introducing Muscle prompts (facial action units) improves total accuracy to 30.36% and F@1 to 19.65%, with notable gains in Hard scenarios (35.71% Acc vs. baseline’s 27.14%). However, Muscle+Context prompts yield mixed results: while Medium accuracy surges to 40.00% (F@1: 33.50%), Easy performance drops to 17.65% Acc, suggesting contextual cues alone may introduce noise without complementary modalities. The integration of Muscle+Context+Body prompts, which combine facial action units (physiological cues), environmental context, and body language, significantly enhances model robustness, achieving 41.07% total accuracy and 31.81% F@1. This configuration demonstrates peak performance in the Easy tier (52.94% Acc, 49.95% F@1), underscoring the synergistic value of multi-modal integration. However, the proposed SoVTP (Ours), which further incorporates others’ emotions into the prompt ensemble, achieves state-of-the-art results: 70.59% Acc (Easy), 48.00% Acc (Medium), and 45.54% total Acc (33.53% F@1). Notably, while SoVTP matches Muscle+Context+Body in Hard scenarios (38.57% Acc), its F@1 slightly declines (28.25% vs. 28.50%). This marginal dip reflects the inherent challenge of the Hard tier, where videos contain sparse or ambiguous social cues (e.g., limited visible faces or interactions). Introducing others’ emotions in such scenarios risks introducing noise due to insufficient emotional cues, slightly reducing precision. Nevertheless, SoVTP’s holistic design—integrating physiological, environmental, behavioral, and social-emotional cues—achieves superior generalization overall, validating its efficacy in balancing specificity and robustness across diverse emotion recognition tasks. The study highlights the critical role of progressive integration: each added modality (action units → context → body language → others’ emotions) systematically enhances model robustness.
This Fig 9 delineates a progressive integration of multi-modal cues for emotion recognition, comparing four prompting strategies—Muscle, Muscle+Context, Muscle+Context+Body, and SoVTP—applied to a masked individual expressing happiness. The Muscle prompt isolates facial action units (AUs): AU06 (cheek raise), AU12 (lip corner pull-up), and AU25 (lip parting), directly associating these physiological markers with the Happy emotion. The Muscle+Context layer augments AU analysis with situational context (e.g., a celebratory setting), reinforcing the Happy classification through environmental alignment. The Muscle+Context+Body tier introduces body language analysis, noting open postures or gestures that correlate with positive affect, further validating the Happy inference. Finally, SoVTP (Ours) synthesizes AU data, contextual signals, body language, and social engagement cues (e.g., gaze direction toward the camera, indicative of attentiveness) to produce a holistic emotion profile. While all tiers converge on Happy, SoVTP distinguishes itself by capturing nuanced behavioral dynamics (e.g., engagement) that refine specificity. For instance, the combined presence of AU06, AU12, and AU25, alongside directed gaze and contextually aligned body language, resolves ambiguities that might arise from isolated AU analysis (e.g., distinguishing genuine happiness from social masking).
This incremental framework underscores the necessity of integrating physiological, environmental, behavioral, and social-interactive cues for robust emotion recognition. The case study exemplifies how SoVTP transcends traditional facial-centric models, addressing limitations in scenarios where emotions are contextually or socially mediated. By systematically layering cues, the framework mitigates false positives and enhances interpretability, demonstrating its methodological superiority in real-world applications requiring nuanced affective computing.
IV-G Qualitative Observations

### v3

@@ IV-C Action units
IV-D Quantitative Results
TABLE IV: Comparison of zero-shot emotion recognition methods across various VLLMs on the DFEW dataset [50], reporting accuracy (%) and F@1 scores. The proposed SoVTP-enhanced Qwen2-VL and LLaVA-NeXT models demonstrate superior performance compared to existing methods.
Methods Backbone Acc (%) F@1
ShareGPT4V [46] CLIP, ViT 25.50 16.85
VILA [47] ViT 24.70 16.22
Video-LLaVA [23] Pre-align ViT 28.60 20.78
Qwen2-VL [45] ViT 31.40 27.63
LLaVA-NeXT [48] SigLIP, ViT 40.80 35.28
Qwen2-VL [45]+SoVTP (Ours) ViT 42.30 42.55
LLaVA-NeXT [48]+SoVTP (Ours) ViT 46.50 46.79
Figure 8: Performance comparison of various VLLMs across seven emotion categories in zero-shot emotion recognition on DFEW dataset [50]. The bar chart depicts the accuracy of VLLMs models including Qwen2-VL+Ours, Qwen2-VL [45], Video-LLAVA [23], VILA [47] and ShareGPT4V [46].
Table IV summarizes the zero-shot emotion recognition performance of several vision-language models on the DFEW dataset, reporting overall accuracy and F@1 scores. Baseline methods such as ShareGPT4V, VILA, Video-LLAVA, Qwen2-VL, and LLAVA-NeXT achieve accuracies ranging from approximately 24.70% to 40.80% and corresponding F@1 scores of 16.22% to 35.28%. By contrast, the SoVTP-enhanced Qwen2-VL and LLAVA-NeXT attain substantially higher results, with Qwen2-VL+SoVTP reaching 42.30% accuracy and 42.55% F@1 and LLAVA-NeXT+SoVTP achieving 46.50% accuracy and 46.79% F@1. Fig. 8 further illustrates these gains at the level of individual emotion categories (Happy, Angry, Sad, Neutral, Surprise, Fear, Disgust), showing that SoVTP-enhanced models consistently outperform their base counterparts across all categories—particularly for Happy, Angry and Surprise—thereby demonstrating the effectiveness of holistic scene prompting in improving zero-shot recognition robustness on a diverse set of emotional expressions.
TABLE V: Comparison of state-of-the-art (SOTA) methods for zero-shot emotion recognition. Previous approaches utilize various types of visual prompts: C: Crop, O: Box, R: Blur Reverse, E: Circle, N: Number, and A: Action Units.
@@ IV-D Quantitative Results
IV-E Visual Prompting
Table V presents a comparative analysis of state-of-the-art (SOTA) methods for zero-shot emotion recognition, evaluating performance across Easy, Medium, and Hard difficulty tiers using Accuracy (Acc) and F1-score (F@1) metrics. It compare the effectiveness of different visual prompting strategies utilized by Qwen2-VL. The proposed SoVTP (Ours) method, which integrates numeric annotations (N), bounding boxes (O), and facial action units (A) as visual prompts, achieves superior performance, significantly outperforming existing approaches. On the Easy tier, SoVTP attains 70.59% Accuracy and 58.00% F@1, a substantial improvement over the second-best method, ReCLIP (35.29% Acc, 24.85% F@1). Similarly, in the Medium tier, SoVTP achieves 48.00% Accuracy and 44.94% F@1, surpassing RedCircle (36.00% Acc, 33.20% F@1) and ReCLIP (36.00% Acc, 32.29% F@1). For the challenging Hard tier, SoVTP maintains robustness with 38.57% Accuracy and 28.25% F@1, outperforming ReCLIP (34.29% Accuracy, 20.75% F@1) and RedCircle (31.43% Acc, 17.71% F@1). In total metrics, SoVTP achieves 45.54% Accuracy and 33.53% F@1, representing a 10.72% absolute Accuracy gain over ReCLIP (34.82% Accuracy, 23.13% F@1) and a 13.40% Accuracy gain over RedCircle (32.14% Accuracy, 19.33% F@1). Notably, the baseline method (Qwen2-VL) with plain text prompts lags significantly across all tiers (Total: 23.21% Accuracy, 10.06% F@1), underscoring the critical role of structured visual prompts. The comparatively lower performance of RedCircle and ReCLIP can be attributed to their reliance on cropping and blurring techniques. This approach increases the model’s focus on specific facial features but sacrifices essential body language and contextual cues, which are vital in more complex situations where emotional cues extend beyond the face. In contrast, the SoVTP approach retains a broader field of view, capturing body language and contextual information. This holistic approach accounts for the significant performance gains observed with SoVTP, especially in challenging conditions, where both facial and contextual cues are crucial for accurate emotion recognition.
@@ IV-E Visual Prompting
TABLE VI: Ablation study for vision and text prompts on Qwen2-VL. Baseline: represents the model’s performance without any additional prompts. Muscle: indicates a visual prompt that uses action units. Muscle+Context: adding extracted context to muscle prompts. Muscle+Context+Body: adding extracted body language to context and muscle prompts. Ours: adding extracted others’ emotions to muscle, context and body language prompts.
@@
IV-F Ablation Study
Table VI presents an ablation study of vision and text prompt components on the Qwen2-VL framework, evaluating the incremental impact of integrating multi-modal cues for zero-shot emotion recognition. The baseline model, devoid of additional prompts, achieves minimal performance (Total: 23.21% Acc, 10.06% F@1), particularly struggling in Easy (11.76% Acc) and Medium (20.00% Acc) tiers. Introducing Muscle prompts (facial action units) improves total accuracy to 30.36% and F@1 to 19.65%, with notable gains in Hard scenarios (35.71% Acc vs. baseline’s 27.14%). However, Muscle+Context prompts yield mixed results: while Medium accuracy surges to 40.00% (F@1: 33.50%), Easy performance drops to 17.65% Acc, suggesting contextual cues alone may introduce noise without complementary modalities. The integration of Muscle+Context+Body prompts, which combine facial action units (physiological cues), environmental context, and body language, significantly enhances model robustness, achieving 41.07% total accuracy and 31.81% F@1. This configuration demonstrates peak performance in the Easy tier (52.94% Acc, 49.95% F@1), underscoring the synergistic value of multi-modal integration. However, the proposed SoVTP (Ours), which further incorporates others’ emotions into the prompt ensemble, achieves state-of-the-art results: 70.59% Acc (Easy), 48.00% Acc (Medium), and 45.54% total Acc (33.53% F@1). Notably, while SoVTP matches Muscle+Context+Body in Hard scenarios (38.57% Acc), its F@1 slightly declines (28.25% vs. 28.50%). This marginal dip reflects the inherent challenge of the Hard tier, where videos contain sparse or ambiguous social cues (e.g., limited visible faces or interactions). Introducing others’ emotions in such scenarios risks introducing noise due to insufficient emotional cues, slightly reducing precision. Nevertheless, SoVTP’s holistic design—integrating physiological, environmental, behavioral, and social-emotional cues—achieves superior generalization overall, validating its efficacy in balancing specificity and robustness across diverse emotion recognition tasks. The study highlights the critical role of progressive integration: each added modality (action units → context → body language → others’ emotions) systematically enhances model robustness.
This Fig 11 delineates a progressive integration of multi-modal cues for emotion recognition, comparing four prompting strategies—Muscle, Muscle+Context, Muscle+Context+Body, and SoVTP—applied to a masked individual expressing happiness. The Muscle prompt isolates facial action units (AUs): AU06 (cheek raise), AU12 (lip corner pull-up), and AU25 (lip parting), directly associating these physiological markers with the Happy emotion. The Muscle+Context layer augments AU analysis with situational context (e.g., a celebratory setting), reinforcing the Happy classification through environmental alignment. The Muscle+Context+Body tier introduces body language analysis, noting open postures or gestures that correlate with positive affect, further validating the Happy inference.
TABLE VII: Ablation study comparing accuracy (%) and inference time (seconds) for different vision and text prompt configurations on the Qwen2-VL model [45]. The results highlight the trade-offs between computational efficiency and performance, with SoVTP (Ours) achieving the highest accuracy and F@1 scores.
Prompts #Params #Time (s) Acc (%) F@1
Baseline [45] 7B 1.44 23.31 10.06
Muscle 7B 2.39 30.36 19.65
Muscle+Context 7B 2.54 33.93 24.51
Muscle+Context+Body 7B 3.15 41.07 31.81
SoVTP (Ours) 7B 3.97 45.54 33.53
TABLE VIII: Performance comparison of Qwen2-VL [45] combined with SoVTP prompting under varying numbers of faces $N$ and different face overlap thresholds $\epsilon$. The metrics reported include accuracy (Acc %) and F@1 score, demonstrating the model’s robustness across different crowd densities and overlap settings. Larger overlap thresholds generally result in decreased performance.
Prompts #Faces $\epsilon$ Acc (%) F@1
Qwen2-VL [45]+ SoVTP $N>6$ 0.0 70.59 58.00
0.2 64.71 51.88
0.4 64.71 51.88
Qwen2-VL [45]+ SoVTP $63<N<=$ 6 0.0 48.00 44.94
0.2 40.00 36.93
0.4 36.00 34.15
Qwen2-VL [45]+ SoVTP $N<=3$ 0.0 38.57 28.25
0.2 37.14 20.33
0.4 37.14 20.33
Figure 12: Ablation study of key hyperparameters. We evaluate the impact of (a) different pixel size ranges ((90,105) through (1440,1680)), (b) maximum output token sizes (128 to 2048) on the chosen performance metric.
Finally, SoVTP (Ours) synthesizes AU data, contextual signals, body language, and social engagement cues (e.g., gaze direction toward the camera, indicative of attentiveness) to produce a holistic emotion profile. While all tiers converge on Happy, SoVTP distinguishes itself by capturing nuanced behavioral dynamics (e.g., engagement) that refine specificity. For instance, the combined presence of AU06, AU12, and AU25, alongside directed gaze and contextually aligned body language, resolves ambiguities that might arise from isolated AU analysis (e.g., distinguishing genuine happiness from social masking). This incremental framework underscores the necessity of integrating physiological, environmental, behavioral, and social-interactive cues for robust emotion recognition. The case study exemplifies how SoVTP transcends traditional facial-centric models, addressing limitations in scenarios where emotions are contextually or socially mediated. By systematically layering cues, the framework mitigates false positives and enhances interpretability, demonstrating its methodological superiority in real-world applications requiring nuanced affective computing.
IV-G Inference time
From Table VII, we observe that richer visual and textural prompts progressively increase inference time—from 1.44s for the baseline with plain text prompt to 3.97s for SoVTP—while yielding steadily higher accuracy (23.31% to 45.54%) and F1 scores (10.06% to 33.53%). Initial additions (e.g., Muscle, Muscle+Context+Body) deliver substantial absolute gains relative to added latency, but the final step to SoVTP shows diminishing returns: a modest 4.5 percentage‐point accuracy increase for an extra 0.8s of inference. We believe this trade-off is acceptable in many applications where higher recognition performance is critical. Moreover, our ablation offers intermediate options: if lower latency is required, one may choose Muscle+Context+Body (3.15s, 41.07% accuracy) or Muscle+Context (2.54s, 33.93% accuracy).
IV-H Impacts of key hyperparameters
IV-H1 Impacts of face overlap threshold
Table VIII evaluates the robustness of Qwen2-VL combined with SoVTP prompting across varying crowd densities—categorized by the number of faces ($N>6$, $63<N<=$ 6, and $N<=3$)—and different face overlap thresholds ($\epsilon=0.0,0.2,0.4$). For each density regime, we report accuracy (Acc %) and F@1 scores at increasing overlap tolerances. The results show that performance is highest in high-density settings ($N>6$) with no overlap ($\epsilon=0.0$) and degrades as $\epsilon$ increases, reflecting sensitivity to occlusion and overlapping detections. Medium- and low-density scenarios exhibit lower absolute metrics overall and similarly decline under larger overlap thresholds. These findings highlight both the benefits of richer others’ emotional cues when many faces are present, illustrating SoVTP’s behavior under different crowd and occlusion conditions.
@@ IV-H1 Impacts of face overlap threshold
IV-H2 Ablation study of pixel size and maximum token
Fig. 12 presents a systematic ablation of two critical SoVTP hyperparameters on Qwen2-VL’s zero‐shot emotion recognition accuracy. In panel (a), varying the maximum pixel size from (90,105) to (1440,1680) yields a sharp rise in accuracy—from approximately 44.6% at the smallest pixel size to 45.54% at (180,210)—after which performance plateaus around 45.54%, indicating diminishing returns for very large pixel‐group ranges. Panel (b) shows that increasing the maximum output token limit progressively improves accuracy, climbing from 34.8% at 128 tokens to a peak of 45.5% at 1024 tokens, and then declining slightly to 42.8% at 2048 tokens, suggesting an optimal token budget of around 1024.
IV-I Qualitative Observations
@@
Figure 14: Visualize the SoVTP-enhanced Qwen2-VL generated samples for different expressions: the left side shows a target individual highlighted by a mask, with the ground-truth emotion label displayed in the top-left corner; the right side shows the generated chain-of-thought reasoning and answer. Yellow text indicates observed emotions of other persons in the scene; green highlights specific facial action units; blue denotes body language cues; and purple captures contextual information. These examples demonstrate how integrating others’ emotional states, fine-grained facial movements, posture and gesture, and situational context informs the model’s final emotion inference.
IV-J Limitations
In Fig. 14, the SoVTP method succeeds in most examples by leveraging multi-person emotional states, action units, body posture, and contextual cues—for example, correctly identifying ‘Fear’ in a chaotic scene with visible widened eyes and tense posture, or ‘Happy’ in a group setting with open eyes and raised mouth corners—demonstrating its holistic reasoning. However, in subfigures (a) and (f), the target faces are blurred or occluded, providing limited context and preventing reliable extraction of facial details and action units; this lack of critical fine-grained cues leads to incorrect emotion labels. Future work could incorporate mechanisms to recover or compensate for missing facial details, such as applying generative inpainting models to approximate occluded regions or using deblurring techniques to enhance video quality when fine-grained cues are unreliable.